OpenStudy (anonymous):

WILL FAN AND MEDAL IF HELP Which ordered pairs are solutions to the equation? y = 3x + 1 A. (6, 18) and (0, 1) B. (6, 19) and (0, 0) C. (6, 19) and (0, 1) D. (19, 6) and (0, 1)

OpenStudy (butterflydreamer):

Substitute in the values into the equation and check to make sure it's correct. For example, if we substitute in the point (6, 18) , we know x = 6 , y = 18 Sub it into your formula: y = 3x + 1 18 = 3*6 +1 18 = 18 + 1 18 = 19 which is INCORRECT .Therefore (6,18) is not a solution. Another example is the point (0, 1), x = 0, y = 1 Sub it into y = 3x+1 and we get 1 = 3*0 +1 1 = 0+1 1 = 1 which is CORRECT , therefore (0,1) is a solution. Basically we want to get it so that it solves the question and the left side of the "=" equals to the right side of the "=" sign.

OpenStudy (anonymous):

thx @butterflydreamer

OpenStudy (butterflydreamer):

you're welcome :)

OpenStudy (anonymous):

i gave u a medal as promised